### v1

@@ 4.1 Encrypted Traffic Proportion Analysis
Common traffic analysis tools, such as Wireshark, can easily distinguish between encrypted protocols like TLS and QUIC, but cannot clearly classify customized encryption protocols. Therefore, we adapted the method proposed by Ren et al. [5] to determine whether traffic is encrypted by calculating the payload entropy of the packets. First, for HTTP protocol packets, we classify them into text, compressed, or media based on the context_type field. For SSL protocol, if the payload entropy is greater than 0.8, it is classified as encrypted. For DNS protocol, packets containing the dns_key field are treated as encrypted. For packets that cannot be classified in the first step, we further classify them based on the magic_number of the payload. The magic_number refers to a unique identifier or pattern of bytes in a packet that helps to determine its type or format, often used to identify specific file types or data structures. If the packet is identified as compressed or media data, it is categorized as compressed or media. Finally, for the remaining packets, we perform a final classification based on entropy values: packets with entropy greater than 0.9 are classified as encrypted, those with entropy less than 0.4 are classified as text, and those with entropy between 0.4 and 0.9 are classified as unknown. Ultimately, we classify the traffic into encrypted traffic, unencrypted traffic, and unknown traffic.
@@
The heatmaps show that the proportion of unencrypted traffic is generally low, with no device type exceeding 8.5% of unencrypted traffic. This indicates that encrypted communication dominates the data transmission of most devices. However, significant variation exists across device types. For instance, camera, doorbell, and speaker devices exhibit an average encrypted traffic share of over 60%, largely driven by the security and privacy requirements in audio and video data transmission. In contrast, the humidifier and plug devices have lower average encrypted traffic proportions of 36% and 38%, while their unknown traffic proportions are the highest among all devices, at 50% and 49%, respectively. Analysis of traffic entropy reveals that most of these devices’ traffic falls within the entropy range of 0.4-0.7. Upon further inspection, it appears that these devices do not use standard encryption protocols but instead rely on proprietary protocols developed by the manufacturers. These proprietary protocols have unstable entropy. By reverse engineering, an network attacker could readily deduce the syntax and semantics of the manufacturer’s protocols, thereby enabling them to manipulate and monitor users’ data. When considering the lifecycle phases of the device, the proportion of encrypted traffic is relatively higher during the setup and interaction phases than in the other two phases. These phases involve crucial services such as device initialization and configuration, cloud computing, and media data transmission. Therefore, more encrypted traffic in these phases ensures communication security.
@@ 4.2 Encryption Protocol Analysis
Table 3: Distrubution of Encryption Protocols Deployed in Devices.
Protocol TLS 1.2 TLS 1.1 TLS 1.3 SSL SSLv2 Other Protocols
Device Num 58 1 3 5 3 10
Through analysis, we found that TLS 1.2 is widely used for encrypted communication across devices, with 58 devices using TLS 1.2 at least once. However, some devices (e.g., IMOU doorbells, Pisen cameras, and Tenda cameras) continue to rely on insecure SSL or TLS 1.1 protocols, which accounts for 7.6% of the total communication traffic on average. Analyzing the destination IP addresses of these traffic showed that most targets are telecommunications service providers (e.g., China Mobile) or cloud service providers (e.g., Alibaba Cloud). Moreover, not all encrypted communications from these devices use SSL or TLS 1.1. There is a mixed use with other protocols (e.g., TLS 1.2). This may result from manufacturers neglecting to fully deprecate outdated protocols during certain functionality upgrades, leaving some connections still using SSL or TLS 1.1. On the other hand, while TLS 1.3 offers significant improvements in performance and security, only three devices in our dataset use it. This indicates that despite its clear advantages, its adoption remains limited, and manufacturers must make greater efforts to implement the latest security protocols. Additionally, we identified 10 devices (e.g., Aqara Camera, Aqara Plug, Xiaovv Camera) using proprietary protocols developed by the manufacturers. The security of these protocols remains to be further analyzed.
@@ 4.3 TLS Certificate Security Analysis
The TLS 1.2 protocol is the most widely used encryption protocol in our dataset. To further understand its security, this section presents a statistical analysis of the certificate for the 58 devices in the dataset that use the TLS 1.2 protocol. The results show that 43 devices use a public key certificate model for communication. The remaining 15 devices used the Pre-Shared Key (PSK) model without relying on certificates. Given the crucial role digital certificates play in TLS encryption security, we further analyzed the self-signed certificate usage and the encryption algorithms’ security. Additionally, we conducted dynamic tests using the mitmproxy to assess the validity of certificates when devices powered on again.
@@ 4.3.1 Certificate Usage
Overall, the certificate security of domestic devices is generally good, but some insecure certificates were still identified. Table 4 statistics the number of TLS implementations that may be insecure, including insecure signature algorithms, insecure cryptographic algorithms, and self-signed certificates. Among the 43 devices, 7 use insecure algorithms, such as the SHA1 signature algorithm and RSA-1024 cryptographic algorithm, primarily for communication with Aliyun and AWS servers. These algorithms are now considered insecure and vulnerable to attacks. Secondly, approximately 39% of the devices use self-signed certificates in at least one TLS 1.2 connection. These certificates are typically signed by CA authorities like ClickPKI or self-signed by the device manufacturers. Many self-signed certificates have excessively long validity periods, with one Lenovo camera certificate valid until 2120. According to current security standards, the validity period of TLS certificates should not exceed 13 months [14]. As pointed out by Dong et al. [15], the lack of public oversight for long-term vendor-signed certificates poses potential vulnerabilities, as they may not meet the stringent standards of trusted public CAs, thereby increasing the security risks within the consumer IoT ecosystem.
@@
TLS 1.2 implementation Camera Doorbell Plug Hub Light Sensor Speaker
Insecure Signature Algorithms 3 0 1 1 0 0 2
Insecure Cryptographic Algorithms 1 0 0 1 0 0 0
Self-signed Certificates 13 1 0 1 1 1 0
@@ 4.3.2 Certificate Validation
Ideally, devices should check whether the certificates from servers are valid. If a forged certificate is detected, the device should either return an error message or reject the TLS connection. However, if a device ignores certificate anomalies and proceeds to establish a connection, the certificate validation mechanism may have vulnerabilities.
To investigate this issue, we first successfully captured traffic during device reconnections and discovered that 35 devices would conduct TLS connection with certificate exchange with 92 servers during the reconnection period. Next, we conducted a mitmproxy certificate replacement attack on these TLS connections. The results are presented in Table 5.
@@
Server handshake failed 1 2
@@
Communicates normally 20 33
The results show that approximately 64% of the servers responded with errors, such as ”Unknown CA,” ”Decrypt error,” ”Close notify” alerts, or actively terminated the connection. Meanwhile, 33% of the servers associated with 20 devices did not verify the certificates. By decrypting the traffic of these servers, we were able to obtain sensitive information such as device logs, configuration details, and authentication credentials. This indicates that many consumer IoT devices follow insecure certificate validation practices, leaving them vulnerable to certificate replacement attacks.
@@ 4.4 PII Exposure in Unencrypted Traffic
The analysis reveals that the exposure of PII in unencrypted communication traffic is minimal in traffic. Only a small amount of PII (including the names of five well-known singers) is detected in the traffic from four speaker devices, and no inputted personal information is found. These singer names appeared in the song requests during the interaction phase. Although this information does not directly reveal the user’s identity, it could still be exploited by malicious third parties to infer user preferences or deliver targeted ads, thereby posing a risk of privacy leakage and data misuse.
@@ 5 Comparisons on Different Regions
This section compares consumer IoT device traffic measurement results in China with other regions. We choose Ren et al.[5] and Paracha M. T. et al.[13] for comparison, as their studies align with our research purpose, focusing on traffic destinations and encryption analysis.
@@ 5.1 Traffic Destinations
In terms of traffic destinations, compared to consumer IoT devices in the United States and the United Kingdom, Chinese devices predominantly communicate domestically. According to Ren et al. [5], the 35 devices in the United Kingdom communicated with 7 overseas regions. Approximately 70% of the United Kingdom device traffic flowed outside their regions (primarily to the United States). In contrast, although Chinese consumer IoT devices contacted 18 overseas countries, overseas traffic only accounted for 3% of total bytes. This highlights Chinese consumer IoT devices’ stronger reliance on domestic services than their United Kingdom counterparts.
@@ 5.2 Encryption Analysis
Encryption practices vary regionally, but unencrypted traffic proportions remain low overall. Ren et al.’s. [5] proposed entropy-based encryption detection, reporting unencrypted traffic proportions below 13% for devices of the United States and the United Kingdom devices. Using similar methods, we come to a similar conclusion: the percentage of unencrypted traffic of consumer IoT devices in China is low, with a peak of 8.5%.
Regarding encryption protocol usage, TLS 1.2 remains the most widely adopted, while TLS 1.3 adoption is still limited. Paracha M. T. et al. [13] analyzed 40 consumer IoT devices in the United States in 2021 and found that TLS 1.2 is the dominant protocol, with minimal use of TLS 1.3 and deprecated versions. Our findings align with these results. Among 70 Chinese devices, 82% utilized TLS 1.2, while 7% still relied on insecure protocols (TLS 1.1/SSL), and 14% employed proprietary protocols developed by manufacturers. For certificate validation, Paracha M. T. et al. [13] examined 32 devices and found that 11 were vulnerable to certificate replacement. In our study, 15 out of 35 devices generated error messages when certificate replaced, while 20 devices exhibited no errors or anomalies, indicating that certificate verification remains a security concern.
@@ 6 Related Work
CN-CIoT CN SC 8 70 2024.3 2.7GB 3 months
Table 6: The Summary of IoT Device Datasets.
@@ 7 Conclusion
We collected 97.5 hours of traffic based on 70 consumer IoT devices across 8 categories and 36 brands. Through analysis, we draw the following key conclusions. First, 99.8% of traffic from Chinese consumer IoT devices is routed domestically, with only 0.2% directed overseas. In contrast, approximately 70% of the United Kingdom device traffic flowed overseas regions. This indicates minimal reliance on foreign services. However, a notable observation is that most devices depend on services from major corporations like Alibaba Cloud, Xiaomi, and Tencent, indicating that most devices rely on the services of these organizations, and these organizations may also obtain more user information from devices. Regarding encryption practices, Chinese consumer IoT devices demonstrate generally robust performance. The unencrypted traffic of Chinese devices peaked at 8.5%, lower than the other regions, and Chinese devices have less PII exposure in unencrypted traffic. Nevertheless, some devices still employ deprecated encryption protocols, insecure TLS certificate algorithms (e.g., SHA1, RSA-1024), and self-signed certificates. More critically, during device certificate replacement, certain devices exhibited an absence of certificate validation mechanisms, exposing communication to potential security and privacy risks.
We open-source our traffic collection guidance and make our dataset publicly available on website [^2] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.

### v2

@@ 4.1 Encrypted Traffic Proportion Analysis
Common traffic analysis tools, such as Wireshark, can easily distinguish between encrypted protocols like TLS and QUIC, but cannot clearly classify customized encryption protocols. Therefore, we adapted the method proposed by Ren et al. [5] to determine whether traffic is encrypted by calculating the payload entropy of the packets. First, for HTTP protocol packets, we classify them into text, compressed, or media based on the context_type field. For the SSL protocol, if the payload entropy is greater than 0.8, it is classified as encrypted. For the DNS protocol, packets containing the dns_key field are treated as encrypted. For packets that cannot be classified in the first step, we further classify them based on the magic_number of the payload. The magic_number refers to a unique identifier or pattern of bytes in a packet that helps to determine its type or format, often used to identify specific file types or data structures. If the packet is identified as compressed or media data, it is categorized as compressed or media. Finally, for the remaining packets, we perform a final classification based on entropy values: packets with entropy greater than 0.9 are classified as encrypted, those with entropy less than 0.4 are classified as text, and those with entropy between 0.4 and 0.9 are classified as unknown. Ultimately, we classify the traffic into encrypted traffic, unencrypted traffic, and unknown traffic.
@@
The heatmaps show that the proportion of unencrypted traffic is generally low, with no device type exceeding 10% of unencrypted traffic during the full device lifecycle. This indicates that encrypted communication dominates the data transmission of most devices. However, significant variation exists across device types. For instance, camera, doorbell, speaker, and TV devices exhibit an average encrypted traffic share of over 60%, largely driven by the security and privacy requirements in audio and video data transmission. In contrast, the humidifier, pet feeder, and plug devices have lower average encrypted traffic proportions of 36%, 34%, and 38%, while their unknown traffic proportions are the highest among all devices, at 50%, 54%, and 49%, respectively. Analysis of traffic entropy reveals that most of these devices’ traffic falls within the entropy range of 0.4-0.7. Upon further inspection, it appears that these devices do not use standard encryption protocols but instead rely on proprietary protocols developed by the manufacturers. These proprietary protocols have unstable entropy. By reverse engineering, we discovered that some devices can deduce the syntax and semantics of the manufacturer’s protocols. Therefore, we suggest that manufacturers can adopt standardized encryption protocols, such as TLS 1.3. In cases where proprietary protocols must be used, they can be combined with packet padding and traffic obfuscation techniques [36] to improve resistance against protocol reverse engineering and effectively reduce the risks of data leakage.
When considering the lifecycle phases of the device, the proportion of encrypted traffic is relatively higher during the setup and interaction phases than in the other two phases. These phases involve crucial services such as device initialization and configuration, cloud computing, and media data transmission. Therefore, more encrypted traffic in these phases ensures communication security.
@@ 4.2 Encryption Protocol Analysis
Table 3: Distribution of Encryption Protocols Deployed in Devices.
Protocol TLS 1.1 TLS 1.2 TLS 1.3 SSL SSLv2 SSLv3 Other Protocols
Setup 1 58 1 5 3 8 11
Idle 2 59 1 2 0 1 11
Interaction 4 54 3 18 14 12 11
Delete 0 48 0 1 0 2 11
Full lifecycle 5 62 5 22 16 17 11
Through analysis, we found that TLS 1.2 is widely used for encrypted communication across devices, with 62 devices using TLS 1.2 at least once. However, some devices (e.g., IMOU doorbells, Pisen cameras, and Tenda cameras) continue to rely on insecure SSL or TLS 1.1 protocols, which accounts for 7.6% of the total communication traffic on average. The devices generally use these protocols during the setup phase and the interaction phase. The setup and interaction phases often involve the transmission of sensitive data such as user identity information, remote control commands, and status update information. If weak encryption protocols are used during these phases, attackers may exploit them through MITM attacks to intercept and reconstruct the transmitted data. Currently, protocols below TLS 1.2 have been deemed insecure by multiple international standards organizations. Devices that continue to use such protocols may fail to comply with data protection requirements such as the General Data Protection Regulation (GDPR) and China’s Cybersecurity Law. Therefore, device manufacturers should review and phase out insecure protocols such as SSL, TLS 1.0, and TLS 1.1, and adopt TLS 1.2 or higher as the standard for secure communication.
Analyzing the destination IP addresses of this traffic showed that most targets are telecommunications service providers (e.g., China Mobile) or cloud service providers (e.g., Alibaba Cloud). Moreover, not all encrypted communications from these devices use SSL or TLS 1.1. There is a mixed use with other protocols (e.g., TLS 1.2). This may result from manufacturers neglecting to fully deprecate outdated protocols during certain functionality upgrades, leaving some connections still using SSL or TLS 1.1. On the other hand, while TLS 1.3 offers significant improvements in performance and security, only three devices in our dataset use it. This indicates that despite its clear advantages, its adoption remains limited, and manufacturers must make greater efforts to implement the latest security protocols. Additionally, we identified 10 devices (e.g., Aqara Camera, Aqara Plug, Xiaovv Camera) using proprietary protocols developed by the manufacturers. The security of these protocols remains to be further analyzed.
@@ 4.3 TLS Certificate Security Analysis
The TLS 1.2 protocol is the most widely used encryption protocol in our dataset. To further understand its security, this section presents a statistical analysis of the certificate for the 63 devices in the dataset that use the TLS 1.2 protocol. The results show that 48 devices use a public key certificate model for communication. The remaining 15 devices used the Pre-Shared Key (PSK) model without relying on certificates. Given the crucial role digital certificates play in TLS encryption security, we further analyzed the self-signed certificate usage and the encryption algorithms’ security. Additionally, we conducted dynamic tests using the mitmproxy to assess the validity of certificates when devices powered on again.
@@ 4.3.1 Certificate Usage
Overall, the certificate security of domestic devices is generally good, but some insecure certificates were still identified. Table 4 statistics the number of TLS implementations that may be insecure, including insecure signature algorithms, insecure cryptographic algorithms, and self-signed certificates. Among the 48 devices, 2 use insecure algorithms, such as the SHA1 signature algorithm and the RSA-1024 cryptographic algorithm, primarily for communication with Aliyun and AWS servers. These algorithms are now considered insecure and vulnerable to attacks. Therefore, device manufacturers can regularly audit the cypher suites and certificate configurations used in their products, eliminate algorithms that are widely recognized as insecure, and ensure that their TLS implementations comply with current industry security standards.
Secondly, approximately 37.5% of the devices use self-signed certificates in at least one TLS 1.2 connection. These certificates are typically signed by CA authorities like ClickPKI or self-signed by the device manufacturers. Many self-signed certificates have excessively long validity periods, with one Lenovo camera certificate valid until 2120. According to current security standards, the validity period of TLS certificates should not exceed 13 months [14]. As pointed out by Dong et al. [15], the lack of public oversight for long-term vendor-signed certificates poses potential vulnerabilities, as they may not meet the stringent standards of trusted public CAs, thereby increasing the security risks within the consumer IoT ecosystem.
@@
TLS 1.2 implementation Camera Doorbell Plug Hub Sensor Speaker Other devices
SU ID IR ID IR DE ID ID ID SU SU DE
Insecure Signature Algorithms 1 0 0 0 0 0 0 1 0 0 0 0
Insecure Cryptographic Algorithms 1 0 0 0 0 0 0 1 0 0 0 0
Self-signed Certificates 10 3 2 1 1 1 1 1 1 1 2 1
@@ 4.3.2 Certificate Validation
Ideally, devices should check whether the certificates from the servers are valid. If a forged certificate is detected, the device should either return an error message or reject the TLS connection. However, if a device ignores certificate anomalies and proceeds to establish a connection, the certificate validation mechanism may have vulnerabilities.
To investigate this issue, we first successfully captured traffic during device reconnections and discovered that 40 devices would conduct TLS connections with certificate exchange with 106 servers during the reconnection period. Next, we conducted a mitmproxy certificate replacement attack on these TLS connections. The results are presented in Table 5.
@@
Server handshake failed 2 3
@@
Communicates normally 23 38
The results show that approximately 65% of the servers responded with errors, such as ”Unknown CA,” ”Decrypt error,” ”Close notify” alerts, or actively terminated the connection, and 35% of the servers associated with 23 devices did not verify the certificates. We decrypted the traffic between the devices and servers and extracted APIs used by the devices, revealing critical device information. For example, the API “/S_PROD1_MEIDIMINI/…/215413.log” exposes device log data; the API “/v3/am/query_device_bind_status” reveals device binding status, and in this API, the fields “device_name” and “device_sk” are exposed. We suspect that “device_sk” may be related to the device’s secret key, which is typically used to verify the identity and ensure the security of the device. These APIs expose sensitive information, if exploited by attackers, they could lead to serious consequences such as log data leakage, unauthorized remote control of the device, or tampering with its binding status. These findings suggest that many consumer IoT devices adopt insecure certificate validation mechanisms, making them vulnerable to certificate replacement attacks. We have reported this issue to the respective device manufacturers and submitted the vulnerability to the China National Vulnerability Database (CNVD).
@@ 4.4 PII Exposure in Unencrypted Traffic
The analysis reveals that the exposure of PII in unencrypted communication traffic is minimal. Only a small amount of PII (including the names of five well-known singers) is detected in the traffic from four speaker devices, and no inputted personal information is found. These singer names appeared in the song requests during the interaction phase. Although this information does not directly reveal the user’s identity, it could still be exploited by malicious third parties to infer user preferences or deliver targeted ads, thereby posing a risk of privacy leakage and data misuse.
@@ 5 Comparisons on Different Regions
This section compares consumer IoT device traffic measurement results in China with those in other regions. We choose Ren et al.[5] and Paracha M. T. et al.[13] for comparison, as their studies align with our research purpose, focusing on traffic destinations and encryption analysis.
@@ 5.1 Traffic Destinations
We have summarized the differences of traffic destinations between China and other regions in the table 6. In terms of traffic destinations, compared to consumer IoT devices in the United States and the United Kingdom, Chinese devices predominantly communicate domestically.
Table 6: Table 7: Comparison of traffic destinations with other regions.
Distribution of Traffic Destinations1 Commonly Contacted Organizations
destination China (77) United States (46) United Kingdom (35) China (77) United States (46) United Kingdom (35)
China 97.00% 2.40% 1.69% Google Amazon Amazon
United States 2.70% 90.00% 72.90% Alibaba Cloud Google Google
Europe 0.20% 7.50% 22.03% Xiaomi Akamai Akamai
Other Reigions 0.10% 0.10% 3.38% Greatbit Microsoft Microsoft
Service Contacted by Devices Baidu Netflix Alibaba
party China (77) United States (46) United Kingdom( 35) Tencent Kingsoft Kingsoft
Support Party 2.68 2.13 2.48 Huawei 21Vianet 21Vianet
Third Party 1.71 0.15 0.14 Tuya Alibaba BeijingHuaxiay
  • 1
    The ratio between the United States and the United Kingdom was estimated in Ren et al.’s work.
According to Ren et al. [5], the 35 devices in the United Kingdom communicated with 7 overseas regions. Approximately 70% of the United Kingdom’s device traffic flowed outside their regions (primarily to the United States). In contrast, although Chinese consumer IoT devices contacted 18 overseas countries, overseas traffic only accounted for 3% of total bytes. This highlights Chinese consumer IoT devices’ stronger reliance on domestic services than their United Kingdom counterparts. The devices in the United Kingdom partially rely on overseas cloud platforms and service providers (such as AWS, Google Cloud, and Akamai), resulting in frequent cross-border transmission of their data to overseas data centres. This increases the risks of data leakage. In contrast, most data transmission activities of Chinese consumer IoT devices occur within the country, indicating a high degree of local service deployment. This phenomenon reflects the high degree of reliance of Chinese manufacturers on local services.
@@ 5.2 Encryption Analysis
Encryption practices vary regionally, but unencrypted traffic proportions remain low overall. Ren et al. [5] proposed entropy-based encryption detection, reporting unencrypted traffic proportions below 13% for devices in the United States and the United Kingdom. Using similar methods, we come to a similar conclusion: the percentage of unencrypted traffic of consumer IoT devices in China is low, with a peak of 10%.
Regarding encryption protocol usage, TLS 1.2 remains the most widely adopted, while TLS 1.3 adoption is still limited. Paracha M. T. et al. [13] analyzed 40 consumer IoT devices in the United States in 2021 and found that TLS 1.2 is the dominant protocol, with minimal use of TLS 1.3 and deprecated versions. Our findings align with these results. Among 70 Chinese devices, 82% utilized TLS 1.2, while 7% still relied on insecure protocols (TLS 1.1/SSL), and 14% employed proprietary protocols developed by manufacturers. For certificate validation, Paracha M. T. et al. [13] examined 32 devices and found that 11 were vulnerable to certificate replacement. In our study, 17 out of 40 devices generated error messages when the certificate was replaced, while 23 devices exhibited no errors or anomalies, indicating that certificate verification remains a security concern.
@@ 6 Related Work
CN-CIoT CN SC 12 77 2024.3 3GB 3 months
@@ 7 Conclusion
We collected 108 hours of traffic based on 77 consumer IoT devices across 12 categories and 38 brands. Through analysis, we draw the following key conclusions. First, 99.75% of traffic from Chinese consumer IoT devices is routed domestically, with only 0.25% directed overseas. In contrast, approximately 70% of the United Kingdom’s device traffic flowed overseas. This indicates minimal reliance on foreign services. However, a notable observation is that most devices depend on services from major corporations like Alibaba Cloud, Xiaomi, and Tencent, indicating that most devices rely on the services of these organizations, and these organizations may also obtain more user information from devices. Regarding encryption practices, Chinese consumer IoT devices demonstrate generally robust performance. The unencrypted traffic of Chinese devices peaked at 10%, lower than the other regions, and Chinese devices have less PII exposure in unencrypted traffic. Nevertheless, some devices still employ deprecated encryption protocols, insecure TLS certificate algorithms (e.g., SHA1, RSA-1024), and self-signed certificates. More critically, during device certificate replacement, certain devices exhibited an absence of certificate validation mechanisms, exposing communication to potential security and privacy risks.
We open-source our traffic collection guidance and make our dataset publicly available on the website [^2] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.